40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 19 18:41:42 2022, yuta, Update, BHD, Contrast measurements for Michelson and ITM-LO ContrastMeasurements.pdf
    Reply  Wed Jul 20 18:07:52 2022, Paco, Update, BHD, BHD MICH test BHD_MICH_OSC.pdf
       Reply  Tue Jul 26 20:54:08 2022, Paco, Update, BHD, BHD MICH test - LO phase control 20220726_BICHD.pdf
          Reply  Wed Jul 27 18:30:50 2022, yuta, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected LOBeamAtBHD.JPGITMXSingleBounceAtBHD.jpg
             Reply  Thu Jul 28 14:34:40 2022, Yehonathan, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected 
                Reply  Thu Jul 28 16:51:55 2022, Tega, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs Screenshot_2022-07-28_16-58-34_LOandASShaking.pngScreenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
                   Reply  Thu Jul 28 20:16:26 2022, Anchal, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs 
                   Reply  Fri Jul 29 18:24:53 2022, Tega, Update, BHD, LO beam power improved by factor of 6 after LO and AS beam alignment LO_and_oldAS_settings.pngBHD_fringe_settings.pngScreenshot_2022-07-29_19-37-39.pngFromTheLeft-AS-POP-LO.JPG
                      Reply  Wed Aug 3 16:00:51 2022, yuta, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
                         Reply  Tue Aug 9 15:33:12 2022, yuta, Update, BHD, BHD fringe contrast improved from 43% to 74% HPC-DCPD_B_OUT_1344118517_ITMY-LO.pngHPC-DCPD_A_OUT_1344118517_ITMY-LO.pngHPC-DCPD_B_OUT_1344118318_ITMX-LO.pngHPC-DCPD_A_OUT_1344118318_ITMX-LO.png
                            Reply  Tue Aug 9 15:50:22 2022, Koji, Update, BHD, BHD fringe contrast improved from 43% to 74% 
                               Reply  Wed Nov 16 15:09:08 2022, yuta, Update, BHD, BHD fringe contrast measured with unwhitening filters ContrastMeasurements20221116_edited.pdf
                         Reply  Tue Oct 25 08:25:00 2022, JC, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping 
                         Reply  Thu Nov 17 11:24:39 2022, JC, HowTo, LSC, Locking MICH 
                            Reply  Tue Nov 29 11:38:37 2022, JC, HowTo, LSC, Lock Single Arm After MICH lock Screen_Shot_2022-11-29_at_11.01.56_AM.png
Message ID: 17044     Entry time: Thu Jul 28 16:51:55 2022     In reply to: 17042     Reply to this: 17045   17046
Author: Tega 
Type: Update 
Category: BHD 
Subject: Shaking test for LO beam AS beam to BHD DCPDs 

[Yuta, Tega, Yehonathan]

To investigate the BHD power imbalance and clipping issues, we did some shaking test of the mirrors in the LO path and AS paths. The results suggests the following:

  • The clipping is happening after the BHD BS in DCPD_A path, as opposed to our initial guess of BHD BS transmission clipping in elog 17040.
  • The LO beam we are seeing is probably a ghost beam from PR2

We performed both PIT and YAW shaking of all mirrors and looked at the output at DCPD_A and DCPD_B, see table below for details. Since we only see the dithering signal in DCPD_A, it suggests that the clipping is ocurring after the BHD BS and is also confined to the path between BHD BS and DCPD_A. We also swapped the camera location from DCPD_A to DCPD_B on ITMY table and confirmed that the beam was clipped for DCPD_A but not for DCPD_B.

This result discounts the possiblity of clipping being responsible for the power imbalance and therefore suggests that the power imbalance may actually be due to BHD BS not being 50:50. From the measurement in elog 17040, the transmission of BHD BS is 44\pm0.3% and the reflectivity is 56\pm0.3%. Note that DCPD_A is the transmission of BHD BS for AS beam, whereas DCPD_B is the reflection of BHD BS for AS beam, elog 16932.

We expect the shaking of PR2 to give no signal in either DCPD_A or DCPD_B when the LO beam is purely in trasmission, however, we see a signal in DCPD_A sugesting that the LO beam transit path through PR2 may not be as expected, i.e. the beam might be exiting the side of PR2 instead of the AR coated surface.

Finally, we measured the coherence between the dithering dof and DCPD_A/B & POP, see attachment 2, where we noticed that both DCPD_A/B have high coherence in the 1Hz-10Hz frequency band whereas ther was no coherence in POP as expected. This suggests that there may also be some small clipping in DCPD_B path.

 

LO Beam Shaking (LO1, LO2, PR2):

color OPTIC DOF Freq OSC Amp (cnts) comment
Black       0 Reference
Blue LO1 YAW 304.4 2000

Signal in DCPD_A & No signal in DCPD_B

Orange LO1 PIT 304.4 2000 No signal
Magenta LO2 YAW 312.2 10000 No signal
Purple LO2 PIT 312.2 10000 Signal in DCPD_A & No signal in DCPD_B
Green PR2 YAW 308.8 20000 Signal in DCPD_A & No signal in DCPD_B
Red PR2 PIT 308.8 20000 Signal in DCPD_A & No signal in DCPD_B

 

AS Beam Shaking (AS1 and AS4)

color OPTIC DOF Freq OSC Amp (cnts) comment
Black       0 Reference
Blue AS1 YAW 305.5 2000

Signal in DCPD_A & No signal in DCPD_B

Orange AS1 PIT 305.5 2000 Signal in DCPD_A & No signal in DCPD_B
Magenta AS4 YAW 313.3 2000 Signal in DCPD_A & No signal in DCPD_B
Purple AS4 PIT 313.3 2000 No signal

 

 

Attachment 1: Screenshot_2022-07-28_16-58-34_LOandASShaking.png  168 kB  Uploaded Thu Jul 28 18:54:12 2022  | Show | Hide all | Show all
Attachment 2: Screenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png  324 kB  Uploaded Thu Jul 28 19:52:42 2022  | Hide | Hide all | Show all
Screenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
ELOG V3.1.3-